# New perspectives on the wave equation for a massive scalar field in Schwarzschild spacetime

### v1

@@ Abstract
For a massive scalar field in a fixed Schwarzschild background, the radial wave equation obeyed by Fourier modes is first studied. In addition to exact solutions at the regular singular points, we find an exact local solution in the neighbourhood of the irregular singular point at infinity. Moreover, we obtain and study the associated integral representation of the massive scalar field. Eventually, the technique developed for the irregular singular point is applied to the homogeneous equation associated to the inhomogeneous Zerilli equation for gravitational perturbations in a Schwarzschild background.
@@ I Introduction
Almost sixty years after Einstein’s seminal papers Einstein1 ; Einstein2 on general relativity, the study of even just the scalar wave equation in a fixed curved background which solves the Einstein equations was still in the early stages. In the seventies, for example, Persides published very detailed papers on the radial wave equation in Schwarzschild space-time, but claimed that this is not related to any known differential equation of mathematical physics Persides1 ; Persides2 ; Persides3 . Forty years after Persides’ work, the remarkable Ph.D. thesis in Ref. Wardell was still claiming that no exact solution of the radial wave equation is known in Schwarzschild space-time.
It has been therefore our aim to exploit the well-established tools of ordinary and partial differential equations to achieve both exact solution of the radial wave equation for Fourier modes of a massive scalar field, and the integral representation of the full scalar field in Schwarzschild geometry. The scalar field case is already of rich physical interest, in light of modern investigations of null infinity Schmidt ; Kehrberger and of its role in studying Hawking radiation Hawking for quantum fields in curved space-time Figari ; Witten .
The paper is organized as follows. Section II is devoted to an original synthesis of the radial wave equation and its reduction to normal form. Section III obtains approximate solutions of such an equation in the neighbourhood of the two regular singular points and of the irregular singular point. Section IV displays the relation between radial wave equation and the Heun equation. Local solutions at the regular singular points are obtained in Sec. V. A full local solution of the radial wave equation at infinity is obtained in Sec. VI. The resulting integral representation of the massive scalar field that solves the wave equation is studied analytically and numerically in Sec. VII. Section VIII is devoted to the application of Sec. VI to the investigation of the homogeneous equation associated to the inhomogeneous Zerilli equation for gravitational perturbations of a Schwarzschild background. Concluding remarks are presented in Sec. IX, and relevant details are provided in the Appendices.
@@ II Radial wave equation and its reduction to normal form
for a scalar field $\psi$ of mass $\mu$ in a fixed Schwarzschild background. This background spacetime is the unique spherically symmetric solution of the vacuum Einstein equations in a way made precise by the Birkhoff theorem HE . In standard spherical coordinates $(t,r,\theta,\phi)$ with $G=c=1$ units, on defining $f(r)=1-\frac{2M}{r}$, the space-time metric is diagonal with components
$g_{\mu\nu}={\rm diag}\Bigr{(}-f,\frac{1}{f},r^{2},r^{2}\sin^{2}\theta\Bigr{)}.$ (2)
Since such components are $t$-independent, one can exploit a Fourier representation according to the expansion
@@
$\left[\frac{d^{2}}{dr^{2}}+\frac{2(r-M)}{r^{2}f}\frac{d}{dr}+\frac{1}{f}\left(\frac{\omega^{2}}{f}-\frac{l(l+1)}{r^{2}}-\mu^{2}\right)\right]R_{l\omega\mu}(r)=0.$ (5)
@@
and looking for an $\alpha(r)$ such that the equation obeyed by $\beta_{l\omega\mu}(r)$ has vanishing coefficient of the first derivative. Indeed, upon setting in Eq. (5)
$p(r)=\frac{2(r-M)}{r^{2}f},\;q_{l\omega\mu}(r)=\frac{1}{f}\left(\frac{\omega^{2}}{f}-\frac{l(l+1)}{r^{2}}-\mu^{2}\right),$ (7)
the general method for achieving normal form Esposito yields
@@
We note that $J_{l\omega\mu}(r)$ has a second-order pole at $r=0$ and at $r=2M$, and hence these are regular singular points, whereas the point $r=\infty$ is an irregular singular point Forsyth as we are going to prove. In order to study the nature of the $r=\infty$ point, one defines the new independent variable $\rho=\frac{1}{r}$, which implies for $\beta_{l\omega\mu}(\rho)$ the differential equation
@@
Since $\gamma_{2\omega\mu}(\rho)$ has a fourth-order pole at $\rho=0$, the point $r=\infty$ is indeed an irregular singular point for Eq. (9).
The material in this section is standard, but nevertheless useful to achieve a self-contained presentation of our subsequent analysis.
III Approximate solutions in the neighbourhood of $r=0,2M$ and $\infty$
@@ III.1 Fourier modes as $r\rightarrow 0$
As $r\rightarrow 0$, the form of $J_{l\omega\mu}(r)$ in Eq. (10) implies that Eq. (9) reduces to a form independent of $l$ and $\omega$, i.e.,
@@
On looking for solutions in the form $\beta(r)=r^{\alpha}$, one finds that the square of $\left(\alpha-\frac{1}{2}\right)$ should vanish, and hence there exist two linearly independent integrals giving rise to
@@
One requiring regularity at the origin one finds $C_{2}=0$, and hence
@@
i.e., the Fourier modes are finite and constant at $r=0$.
@@ III.2 Fourier modes as $r\rightarrow 2M$
As $r\rightarrow 2M$, by virtue of Eq. (10) we find that
@@
This implies that Eq. (9) reduces to
@@
where $D_{1}$ and $D_{2}$ are suitable constants.
@@ III.3 Fourier modes at infinity
Last but not least, the potential term $J_{l\omega}$ in Eq. (9) has the following Poincaré asymptotic expansion as $r\rightarrow\infty$:
@@
Thus, as $r\rightarrow\infty$, $\beta_{l\omega\mu}(r)$ is (up to a multiplicative constant) the product of the leading order $\beta_{\omega,\infty}(r)=e^{\pm r\sqrt{\mu^{2}-\omega^{2}}}$ times a function that we do not strictly need here, but is part of an analysis that we will perform in Sec. VI.
Note also that, by virtue of Eq. (4), one has to integrate over all values of $\omega$ in order to evaluate the large-$r$ scalar field. Thus, the Fourier modes at large $r$ have to be considered in the following intervals for the $\omega$ variable:
@@
This remark will be exploited in Sec. VII.
@@ IV Connecting the radial with the Heun equation
The reduction to canonical form of Eq. (5) helped us in searching for the singular points $r_{0}=0$, $r_{2M}=2M$ and for the one at infinity. Now we are going to see that Eq. (5) belongs to the class of confluent Heun equations Ronveaux . In order to prove it, we rely upon the change of independent variable frequently used in the literature, i.e.
@@
In this way, the two finite singularities $r=0,2M$ correspond to $z=1,0$, respectively, and the physical requirement $r\geq 0$ leads to the restriction $z\leq 1$. The equation (5) henceforth takes the form
@@
Thanks to this change of variable, such equation is nothing but a confluent Heun equation in the so-called non-symmetric canonical form (see Appendix A for a general overview about confluent Heun equations). The non-symmetric canonical form is obtained by inserting in the confluent Heun equation
@@
where $H(z)$ is the confluent Heun function.
@@ V Local solutions at the regular singular points
In our specific case, upon defining the five Heun parameters (no confusion should arise with $\alpha$ and $\beta$ of Sec. II)
@@
since $\alpha_{3}=0$ for Eq. (32). Starting from Eq. (35), one can construct the complete Frobenius solution at $z=0$ (corresponding to $r=2M$), that reads (cf. Refs. Fiziev ; Philipp ; Vieira ; Chen ; Minucci )
@@
In order to visualize the results obtained so far, we plot the real (Fig. 1, Fig. 3) and imaginary (Fig. 2, Fig. 4) part of Eqs. (36), (38), with the following choice for the free parameters:
@@
Figure 1: In this plot, the real part of Eq. (36) is shown for $z\in[-0.01,0.01]$. The $s$-wave near $z=0$ (i.e. $r=2M$) is very close to zero inside the black hole and diverges on the event horizon, then it exhibits strong oscillations.
Figure 2: The imaginary part of Eq. (36) is shown for $z\in[-0.01,0.01]$ The phase of the $s$-wave near $z=0$ behaves qualitatively in the same manner of the real part.
Figure 3: The real part of Eq. (38) is shown for $z\in[0.9,1]$. Near $z=1$ (i.e. $r=0$) the $s$-wave exhibits a plateau, then an overall downward trend as $z$ decreases.
Figure 4: The imaginary part of Eq. (38) is shown for $z\in[0.9,1]$. The phase shows qualitatively the same behaviour of the real part but with sligthly higher values.
VI Full local solution for Fourier modes in the neighbourhood of $r=\infty$
When $r\rightarrow\infty$ one can evaluate $\beta_{l\omega\mu}(r)$, and hence the Fourier modes $R_{l\omega\mu}(r)$, by completing the analysis initiated in Eqs. (28)-(30). For this purpose, we recognize from such equations that the differential equation (9) for $\beta_{l\omega\mu}(r)$ can be written in the form
@@ VI Full local solution for Fourier modes in the neighbourhood of $r=\infty$
where $A_{0}=\omega^{2}-\mu^{2}$, while $A_{1}$ and $A_{2}$ have been already defined in Eqs. (29) and (30), respectively, and all subsequent coefficients can be computed from the asymptotic expansion of Eq. (10) at large $r$. At this stage, Eq. (41) suggests looking for $\beta_{l\omega\mu}(r)$ at large $r$ in the form (setting $\varepsilon=\pm 1$ and considering $\zeta$ as a parameter to be determined as shown below)
@@
The insertion of Eq. (42) into Eq. (41) yields
@@
jointly with infinitely many other equations for all subsequent $\gamma_{k}$ coefficients. Since they should all vanish, we obtain linear algebraic equations for $B_{1},B_{2},...,B_{\infty}$. In addition, since we look for a bounded solution at large $r$, we restrict ourselves to the $\varepsilon=-1$ case. For instance, from Eq. (45) we evaluate the explicit form of $\zeta$, while from Eq. (46) we obtain $B_{1}$, i.e.
@@
By virtue of Eq. (6), we find eventually the large-$r$ Fourier modes in the form (cf. Refs. Fiziev ; Philipp ; Minucci )
@@
Note that, by considering only the $k=0,1,2$ terms in Eq. (41), one would obtain the Whittaker equation (a modified form of the confluent hypergeometric equation), solved by the two Whittaker functions. However, all negative powers of $r$ occur in Eq. (41), and hence one needs our factorization in Eq. (49), where the exponential cancels the effect of $A_{0}$, and the $B_{s}$ functions of $l,\omega$ compensate the negative powers of $r$ in the operator provided that a non-integer power of $r$ is included in our factorized ansatz, in agreement with the general method of Poincaré for linear differential equations Poincare ; Poincare2 .
In Appendix C we offer a theoretical motivation for the ansatz given by Eq. (42) and the explicit computation of $\zeta$ and the first $A_{k}$, $B_{s}$ coefficients. In order to visualize such a solution, we consider Eq. (49) with the same choice for parameters given by Eq. (40) (this time we don’t have integration constants like $C_{1}$ and $C_{2}$), ending up with two plots: one for the real part (Fig. 5) of the field and the other for the imaginary one (Fig. 6). The relation with the local results by Persides in the massless case Persides1 is discussed in Appendix D.
Figure 5: The real part of Eq. (49) is plotted with parameters given by Eq. (40) summing over $s$ from $1$ to $5$ for $r\in[10M,15M]$. The oscillations of the field become smaller and smaller as $r$ grows.
Figure 6: The imaginary part of Eq. (49) is plotted with parameters given by Eq. (40) summing over $s$ from $1$ to $5$ for $r\in[10M,15M]$. The phase follows qualitatively the same behaviour of the real part.

### v2

# New Perspectives on the Irregular Singular Point of the Wave Equation for a Massive Scalar Field in Schwarzschild Space-Time

@@ Abstract
For a massive scalar field in a fixed Schwarzschild background, the radial wave equation obeyed by Fourier modes is first studied. After reducing such a radial wave equation to its normal form, we first study approximate solutions in the neighborhood of the origin, horizon and point at infinity, and then we relate the radial with the Heun equation, obtaining local solutions at the regular singular points. Moreover, we obtain the full asymptotic expansion of the local solution in the neighborhood of the irregular singular point at infinity. We also obtain and study the associated integral representation of the massive scalar field. Eventually, the technique developed for the irregular singular point is applied to the homogeneous equation associated with the inhomogeneous Zerilli equation for gravitational perturbations in a Schwarzschild background.
Schwarzschild space-time; black hole physics; radial wave equation; hyperbolic equations on manifolds; Klein–Gordon equation; Heun equation; irregular singular point; Poincaré asymptotic expansion
@@ I Introduction
Almost a hundred and ten years after the creation of general relativity [1, 2], even the study of just the scalar wave equation in a fixed curved background, which solves the Einstein equations, is still in the early stages. In the seventies, for example, very detailed papers were published on the radial wave equation in Schwarzschild space-time, but they claimed that it is not related to any known differential equation of mathematical physics [3, 4, 5]. Forty years after this work, the dissertation in Ref. [6] was still claiming that no exact solution of the radial wave equation is known in Schwarzschild space-time.
Every partial differential equation studied in physics and mathematics is a world of its own, and the choice of the appropriate techniques depends on the assumptions made: the nature of the physical phenomenon, its mathematical formulation, the ambient space-time manifold, the regularity required for the solutions and the existence or lack of symmetries. It has been, therefore, our aim to exploit the well-established tools of ordinary and partial differential equations to obtain both the exact solution of the radial wave equation for Fourier modes of a massive scalar field and the integral representation of the full scalar field in Schwarzschild geometry. The scalar field case is already of considerable physical interest in light of modern investigations of null infinity [7, 8] and of its role in studying Hawking radiation [9] for quantum fields in curved space-time [10, 11]. The spherical symmetry of the fixed space-time geometry makes it possible to exploit Fourier modes, which do not exist in a generic space-time without any symmetry.
The paper is organized as follows. Section II is devoted to an original synthesis of the radial wave equation and its reduction to normal form. Section III obtains approximate solutions of such an equation in the neighborhood of two regular singular points and of an irregular singular point. Section IV gives the radial wave equation in the Heun equation form. Local solutions at the regular singular points are obtained in Section V. A full local solution of the radial wave equation at infinity is obtained in Section VI. The resulting integral representation of the massive scalar field that solves the wave equation is studied analytically and numerically in Section VII. Section VIII is devoted to the application of Section VI to the investigation of the homogeneous equation associated with the inhomogeneous Zerilli equation for gravitational perturbations of a Schwarzschild background. The results obtained and open problems are presented in Section IX, and relevant details are provided in the appendices.
@@ II Radial Wave Equation and Its Reduction to Normal Form
for a scalar field $\psi$ of mass $\mu$ in a fixed Schwarzschild background (thus, we are not studying a coupled system of Einstein and scalar-field equations, nor are we studying a scalar self-force problem). This background space-time is the unique spherically symmetric (This means that space-time admits the group $SO(3)$ as a group of isometries, with the group orbits given by spacelike two-surfaces.) solution of the vacuum Einstein equations in a way made precise by the Birkhoff theorem [12]. In standard spherical coordinates $(t,r,\theta,\phi)$ where $t\in{\bf R}\cup\{-\infty\}\cup\{+\infty\},r\geq 0,\theta\in[0,\pi],\phi\in[0,2\pi]$, with $G=c=1$ units, on defining $f(r)=1-\frac{2M}{r}$, the space-time metric is diagonal:
$g=-f(r)dt^{2}+\frac{dr^{2}}{f(r)}+r^{2}(d\theta^{2}+\sin^{2}\theta d\phi^{2}).$ (2)
Since such components are $t$-independent, one can exploit a Fourier representation according to the expansion [6] ($Y_{lm}$ being the spherical harmonics on the $2$-sphere, and $C_{lm}$ being coefficients of linear combination (The $C_{lm}$ coefficients can be derived by inverse Fourier transform following standard techniques, but we do not need their values in our analysis)).
@@
$\left[\frac{d^{2}}{dr^{2}}+\frac{2(r-M)}{r^{2}f(r)}\frac{d}{dr}+\frac{1}{f(r)}\left(\frac{\omega^{2}}{f(r)}-\frac{l(l+1)}{r^{2}}-\mu^{2}\right)\right]R_{l\omega\mu}(r)=0.$ (5)
@@
and looking for an $\alpha(r)$ such that the equation obeyed by $\beta_{l\omega\mu}(r)$ has vanishing coefficient of the first derivative. Indeed, upon setting in Equation (5)
$p(r)=\frac{2(r-M)}{r^{2}f(r)},\;q_{l\omega\mu}(r)=\frac{1}{f(r)}\left(\frac{\omega^{2}}{f(r)}-\frac{l(l+1)}{r^{2}}-\mu^{2}\right),$ (7)
the general method for achieving normal form [13] yields
@@
We note that $J_{l\omega\mu}(r)$ has a second-order pole at $r=0$ and at $r=2M$, and hence these are regular singular points, whereas the point $r=\infty$ is an irregular singular point [14] as we are going to prove. In order to study the nature of the $r=\infty$ point, one defines the new independent variable $\rho=\frac{1}{r}$, which implies for $\beta_{l\omega\mu}(\rho)$ the differential equation
@@
Since $\gamma_{2\omega\mu}(\rho)$ has a fourth-order pole at $\rho=0$, the point $r=\infty$ is indeed an irregular singular point for Equation (9).
The material in this section is standard but nevertheless useful for achieving a self-contained presentation of our subsequent analysis.
III Approximate Solutions in the Neighborhood of $r=0,2M$ and $\infty$
@@ III.1 Fourier Modes as $r\rightarrow 0$
As $r\rightarrow 0$, the form of $J_{l\omega\mu}(r)$ in Equation (10) implies that Equation (9) reduces to a form independent of $l$ and $\omega$, i.e.,
@@
On looking for solutions in the form $\beta_{0}(r)=r^{\alpha}$, one finds that the square of $\left(\alpha-\frac{1}{2}\right)$ should vanish. Hence there exist two linearly independent integrals giving rise to
@@
On requiring regularity at the origin one finds $C_{2}=0$, and hence
@@
i.e., the Fourier modes are finite and constant at $r=0$. This agrees with the property in the massless case found in Ref.  [3]: there exists one solution finite at the origin, which is physically preferable for this reason.
@@ III.2 Fourier Modes as $r\rightarrow 2M$
As $r\rightarrow 2M$, by virtue of Equation (10) we find that
@@
This implies that Equation (9) reduces to
@@
where $D_{1}$ and $D_{2}$ are suitable constants. Interestingly, this limiting behavior agrees with the massless case [3]: the two solutions, with coefficients $D_{1}$ and $D_{2}$, remain bounded on the horizon $r=2M$, and no solution approaches zero as $r\rightarrow 2M$. These properties are closely related to the possibility of destroying the black hole and to the radiation of higher multipole moments when a small scalar particle falls into the black hole [15].
@@ III.3 Fourier Modes at Infinity
Last but not least, the potential term $J_{l\omega}$ in Equation (9) has the following Poincaré asymptotic expansion as $r\rightarrow\infty$:
@@
Thus, as $r\rightarrow\infty$, $\beta_{l\omega\mu}(r)$ is (up to a multiplicative constant) the product of the leading order $\beta_{\omega,\infty}(r)=e^{\pm r\sqrt{\mu^{2}-\omega^{2}}}$ times a function that we do not strictly need here, but is part of an analysis that we will perform in Section VI.
Note also that, by virtue of Equation (4), one has to integrate overall values of $\omega$ in order to evaluate the large-$r$ scalar field. Thus, the Fourier modes at large $r$ have to be considered in the following intervals for the $\omega$ variable:
@@
This remark will be exploited in Section VII.
@@ IV Connecting the Radial with the Heun Equation
The reduction to canonical form of Equation (5) helped us in searching for the singular points $r_{0}=0$, $r_{2M}=2M$ and for the one at infinity. Now we are going to see that Equation (5) belongs to the class of confluent Heun equations defined in Appendix A [16]. In order to prove it, we rely upon the change of independent variable frequently used in the literature, i.e.,
@@
In this way, the two finite singularities $r=\{0,2M\}$ correspond to $z=\{1,0\}$, respectively, and the physical requirement $r\geq 0$ leads to the restriction $z\leq 1$. The Equation (5) henceforth takes the form
@@
Thanks to this change of variable, such equation is nothing but a confluent Heun equation in the so-called non-symmetric canonical form. Such a form is obtained by exploiting
@@
where $H(z)$ solves the confluent Heun equation.
@@ V Local Solutions at the Regular Singular Points
In our specific case, upon defining the five Heun parameters (no confusion should arise with $\alpha$ and $\beta$ of Section II)
@@
since $\alpha_{3}=0$ for Equation (32). Starting from Equation (35), one can construct the complete Frobenius solution at $z=0$ (corresponding to $r=2M$), that reads (cf. Refs. [17, 18, 19, 20, 21])
@@
In order to visualize the results obtained so far, we plot in Figures 1 and 2 the real and imaginary part of Equation (36), respectively. Similarly, we plot in Figures 3 and 4 the real and imaginary part of Equation (38), respectively. In all these Figures, we make the following choice for the free parameters:
@@
In order to summarize the physical content of our results regarding the $s$-wave, the real part of the local solution slightly inside the black hole is very close to zero, while outside the event horizon, it exhibits strong oscillations, both in amplitude and in frequency. The imaginary part of the same solution follows the same behavior both qualitatively and quantitatively. Moreover, both the real and imaginary parts of the local solution at the origin possess a plateau very close to the z = 1 (i.e., $r=0$) point, and then decrease as z decreases.
Figure 1: In this plot, the real part of Equation (36) is shown for $z\in[-0.01,0.01]$. The $s$-wave near $z=0$ (i.e., $r=2M$) is very close to zero inside the black hole and diverges on the event horizon, and it exhibits strong oscillations.
Figure 2: The imaginary part of Equation (36) is shown for $z\in[-0.01,0.01]$ The phase of the $s$-wave near $z=0$ behaves qualitatively in the same manner as the real part.
Figure 3: The real part of Equation (38) is shown for $z\in[0.9,1]$. Near $z=1$ (i.e., $r=0$) the $s$-wave exhibits a plateau, then an overall downward trend as $z$ decreases.
Figure 4: The imaginary part of Equation (38) is shown for $z\in[0.9,1]$. The phase shows qualitatively the same behavior as the real part but with slightly higher values.
VI Full Local Solution for Fourier Modes in the Neighborhood of $r=\infty$
When $r\rightarrow\infty$ one can evaluate $\beta_{l\omega\mu}(r)$, and hence the Fourier modes $R_{l\omega\mu}(r)$, by completing the analysis initiated in Equations (28)–(30). For this purpose, we recognize from such equations that the differential Equation (9) for $\beta_{l\omega\mu}(r)$ can be written in the form
@@ VI Full Local Solution for Fourier Modes in the Neighborhood of $r=\infty$
where $A_{0}=\omega^{2}-\mu^{2}$, while $A_{1}$ and $A_{2}$ have been already defined in Equations (29) and (30), respectively. All subsequent coefficients can be computed from the asymptotic expansion of Equation (10) at large $r$. At this stage, Equation (41) suggests looking for $\beta_{l\omega\mu}(r)$ at large $r$ in the form (setting $\varepsilon=\pm 1$ and considering $\zeta$ as a parameter to be determined as shown below)
@@
The insertion of Equation (42) into Equation (41) yields
@@
jointly with infinitely many other equations for all subsequent $\gamma_{k}$ coefficients. Since they should all vanish, we obtain linear algebraic equations for $B_{1},B_{2},...,B_{\infty}$. In addition, since we look for a bounded solution at large $r$, we restrict ourselves to the $\varepsilon=-1$ case. (We are here studying a local solution which only holds at large $r$ [3, 5]. Thus, the $\varepsilon=1$ mode cannot be avoided by studying conditions near the horizon $r=2M$.) For instance, from Equation (45) we evaluate the explicit form of $\zeta$, while from Equation (46) we obtain $B_{1}$, i.e.,
@@
By virtue of Equation (6), we find eventually the large-$r$ Fourier modes in the form (cf. Refs. [17, 18, 21])
@@
Note that, by considering only the $k=0,1,2$ terms in Equation (41), one would obtain the Whittaker equation (a modified form of the confluent hypergeometric equation), solved by the two Whittaker functions. However, all negative powers of $r$ occur in Equation (41), and hence one needs our factorization in Equation (49), where the exponential cancels the effect of $A_{0}$, and the $B_{s}$ functions of $l,\omega,\mu$ compensate the negative powers of $r$ in the operator provided that a non-integer power of $r$ is included in our factorized ansatz. This procedure agrees with the general method of Poincaré for linear differential equations [22, 23].
In Appendix C we offer a theoretical motivation for the ansatz given by Equation (42) and the explicit computation of $\zeta$ and the first $A_{k}$, $B_{s}$ coefficients. In order to visualize such a solution, we consider Equation (49) with the same choice for parameters given by Equation (40) (this time we do not have integration constants like $C_{1}$ and $C_{2}$), ending up with two plots: one for the real part (Figure 5) of the field and the other for the imaginary one (Figure 6). The relation with the local results by Persides in the massless case [3] is discussed in Appendix D.
Figure 5: The real part of Equation (49) is plotted with parameters given by Equation (40) summing over $s$ from $1$ to $5$ for $r\in[10M,15M]$. The oscillations of the field become smaller and smaller as $r$ grows.
The plots obtained in our large-$r$ analysis clearly describe the vanishing of the field (both real and imaginary parts) at spacelike infinity due to damped oscillations. This is in agreement with our mathematical ansatz and physical requirements.
Figure 6: The imaginary part of Equation (49) is plotted with parameters given by Equation (40) summing over $s$ from $1$ to $5$ for $r\in[10M,15M]$. The phase follows qualitatively the same behavior of the real part.